# When, Why and How Much? Adaptive Learning Rate Scheduling by Refinement

### v1

@@ 3 Experiments
Table 1: Logistic Regression Experiments (Train Error Rate %).
Problem Stepwise Cosine Linear Refined $\left\|g\right\|_{2}^{2}$ Refined $\left\|g\right\|_{1}$ Refined
Aloi $13.38$$\pm 0.05$ $12.69$$\pm 0.05$ $12.76$$\pm 0.06$ $\mathbf{11.75}$$\pm 0.01$ $12.24$$\pm 0.04$ $12.30$$\pm 0.05$
Glass $31.39$$\pm 0.16$ $30.82$$\pm 0.22$ $30.72$$\pm 0.18$ $\mathbf{30.05}$$\pm 0.44$ $\mathbf{29.62}$$\pm 0.37$ $\mathbf{30}$$\pm 0.41$
Iris $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.39}$$\pm 0.000$ $\mathbf{1.46}$$\pm 0.07$ $\mathbf{1.46}$$\pm 0.07$ $\mathbf{1.46}$$\pm 0.07$
Letter $22.24$$\pm 0.008$ $\mathbf{22.24}$$\pm 0.01$ $\mathbf{22.20}$$\pm 0.02$ $\mathbf{22.23}$$\pm 0.03$ $\mathbf{22.20}$$\pm 0.03$ $\mathbf{22.20}$$\pm 0.03$
Pendigits $4.70$$\pm 0.02$ $4.67$$\pm 0.01$ $\mathbf{4.62}$$\pm 0.03$ $\mathbf{4.56}$$\pm 0.02$ $\mathbf{4.58}$$\pm 0.02$ $\mathbf{4.56}$$\pm 0.04$
Sensorless $11.84$$\pm 0.09$ $11.30$$\pm 0.09$ $11.29$$\pm 0.09$ $10.71$$\pm 0.08$ $\mathbf{10.08}$$\pm 0.05$ $\mathbf{10.11}$$\pm 0.05$
Vehicle $18.83$$\pm 0.09$ $18.49$$\pm 0.05$ $18.55$$\pm 0.06$ $\mathbf{18.21}$$\pm 0.12$ $\mathbf{18.19}$$\pm 0.08$ $\mathbf{18.21}$$\pm 0.1$
Vowel $23.43$$\pm 0.08$ $22.99$$\pm 0.09$ $22.94$$\pm 0.10$ $\mathbf{22.48}$$\pm 0.12$ $\mathbf{22.44}$$\pm 0.08$ $\mathbf{22.41}$$\pm 0.08$
To validate the effectiveness of our method on convex problems, we performed a comparison across 8 commonly used classification problems from the LIBSVM repository, with separable problems excluded (see Section 3.5). We used a logistic regression loss together with the Adam optimizer with $\beta=(0.9,0.95)$, with batch size 16 and trained for 100 epochs. The Refined schedules were calculated by using the linear decay schedule to generate the initial gradient norm sequence. Table 1 demonstrates that both the linear decay schedule and our refinement schedule consistently either match or out-perform the cosine schedule. The linear decay schedule matches the cosine schedule on every problem (up to statistical significance), and out-performs it on two problems. The Refined schedule further out-performs the linear decay schedule, either matching of exceeding it across the board.
@@ 3.2 Large-scale schedule benchmark
We performed a large-scale comparison of schedules across common deep learning benchmarks:
@@
For a small-scale image-classification problem, we chose CIFAR10 (Krizhevsky, 2009). A high-performance Wide ResNet architecture was used (Zagoruyko and Komodakis, 2016). Test error percentage is reported.
@@
A more realistic yet still small scale benchmark. We used a compact DenseNet architecture (Huang et al., 2017) to increase the diversity of model architectures tested. Test error percentage is reported.
@@
We used a ResNet-50 architecture (He et al., 2016) for the ImageNet (Russakovsky et al., 2015) classification task. Note that we used a standard (non-SOTA) data-augmentation pipeline following other recent comparisons in the optimization literature. Test error percentage is reported.
@@
A small-scale German-English translation task (Cettolo et al., 2014) using a LSTM model (Wiseman and Rush, 2016). Test Perplexity is reported.
@@
A modern small (162M parameter) GPT-style auto-regressive transformer model (Radford et al., 2019) trained on the large Book-Wiki corpus (Zhu et al., 2015). Test Perplexity is reported.
@@
A masked autoencoder variant (Liu et al., 2019) also trained on the Book-Wiki corpus. Test Perplexity is reported.
@@
A high-performance ImageNet classifier using a Vision Transformer (Dosovitskiy et al., 2021). In contrast to our ResNet-50 experiments, here we use modern data-augmentation, following standard practice for ViT training. Test error percentage is reported.
@@
A modern recommendation system engine (Naumov et al., 2019) trained on the open Criteo Kaggle Display Advertising dataset. Test Accuracy is reported.
@@
A large stacked U-Net architecture (Sriram et al., 2020) trained on the fastMRI dataset (Zbontar et al., 2018), an image-to-image regression task from the medical imaging domain. Test set metric $100\cdot(1-\text{SSIM})$ is reported.
@@
For each problem, we first performed a sweep of learning rates on a grid [$10^{i}$, $2\times 10^{i}$, $5\times 10^{i}$] for varying $i$, separately for each schedule. We then ran multiple seeds using the best learning-rate from the grid search. Mean and standard error of the mean was tabulated for each schedule. The best result for each method, up to a statistical significance level of 0.05 using a paired two-sample t-test, is highlighted in bold. Specific details of the hyper-parameters used for each problem are given in Appendix G. All non-adaptive schedules included a fixed learning-rate warmup, with length following standard practices for the problem. Our step-wise schedule uses a 30-60-90 percent tenthing, following standard practices from ImageNet training. As in the convex experiments, the Refined schedules were calculated by using the linear schedule to generate the initial gradient norm sequence.
Tables 2 & 3 show the results. We break the schedules into two categories, classical and modern. The modern schedules consistently outperform the classical schedules, often by large margins. Although this is common folk-law in deep learning, we are not aware of any existing large-scale experiments establishing this. Our comparison of modern schedules shows a clear hierarchy among the schedules. The Stepwise schedule is dominated by the Cosine schedule, and the Linear schedule matches or outperforms the Cosine schedule on all problems except ViT. The refined schedule further outperforms the Linear schedule on 5 of the 10 problems, but shows mixed results on ViT and RCNN. The refinement process produced degenerate schedules that fail on the CIFAR problems, we discuss this in Appendix 3.5.
@@
Problem Flat 1/t 1/sqrt Offset 1/t Offset 1/sqrt Linear
CIFAR10 $8.04$$\pm.13$ $5.42$$\pm.28$ $6.37$$\pm.41$ $9.23$$\pm.08$ $6.58$$\pm.06$ $\mathbf{4.35}$.$\pm.05$
CIFAR100 $30.43$$\pm.20$ $26.58$$\pm.11$ $29.09$$\pm.16$ $32.80$$\pm.07$ $27.62$$\pm.13$ $\mathbf{22.11}$.$\pm.08$
ImageNet $33.00$$\pm.14$ $26.48$$\pm.06$ $28.35$$\pm.05$ $47.94$$\pm.08$ $27.34$$\pm.06$ $\mathbf{23.11}$.$\pm.07$
IWSLT14 $8.07$$\pm.02$ $7.62$$\pm.01$ $7.52$$\pm.01$ $12.89$$\pm.06$ $8.48$$\pm.01$ $\mathbf{7.10}$.$\pm.01$
GPT $20.20$$\pm.000$ $18.99$$\pm.04$ $19.48$$\pm.02$ $27.85$$\pm.05$ $22.88$$\pm.006$ $\mathbf{18.60}$.$\pm.02$
RoBERTa $4.52$$\pm.005$ $4.25$$\pm.007$ $4.33$$\pm.01$ $5.33$$\pm.02$ $5.15$$\pm.02$ $\mathbf{3.94}$.$\pm.007$
DLRM $\mathbf{20.95}$.$\pm.006$ $47.59$$\pm 6.45$ $45.99$$\pm 5.98$ $\mathbf{20.94}$.$\pm.007$ $20.99$$\pm.009$ $\mathbf{20.94}$.$\pm.006$
MRI $\mathbf{9.00}$.$\pm.04$ $8.91$$\pm.01$ $\mathbf{8.98}$.$\pm.04$ $9.53$$\pm.08$ $9.16$$\pm.05$ $\mathbf{8.88}$.$\pm.02$
ViT $30.11$$\pm.27$ $28.36$$\pm.40$ $28.53$$\pm.15$ $73.84$$\pm 6.08$ $50.36$$\pm 12.39$ $\mathbf{24.82}$.$\pm.31$
RCNN $65.43$$\pm.12$ $63.38$$\pm.05$ $64.13$$\pm.10$ $79.32$$\pm.07$ $69.25$$\pm.07$ $\mathbf{60.98}$.$\pm.02$
@@
RCNN $61.76$$\pm.06$ $\mathbf{61.00}$.$\pm.04$ $\mathbf{60.98}$.$\pm.02$ $61.88$$\pm.02$
@@ 3.3 Cosine schedule ablations
To further investigate the failure modes of the cosine schedule, we ran a series of shorter duration training runs on CIFAR10. Our premise was that the cosine schedule is heavily over-fit to long training duration computer vision problems. As shown in Table 4, the cosine schedule begins to under-perform the linear decay schedule and the refined schedules when training for less than 30 epochs. In contrast, while the refined schedule also under-performs for longer duration training it has no statistically significant differences from the linear decay schedule for shorter duration training, where they both perform consistently better than the cosine schedule.
Table 4: CIFAR10 Training Time Ablations (Test Error %).
Schedule 1 Epoch 5 Epochs 15 Epochs 30 Epochs
Cosine $35.80$$\pm 0.27$ $15.07$$\pm 0.11$ $7.99$$\pm 0.06$ $\mathbf{6.08}$$\pm 0.03$
Linear Decay $\mathbf{33.84}$$\pm 0.17$ $\mathbf{14.36}$$\pm 0.07$ $\mathbf{7.65}$$\pm 0.07$ $\mathbf{6.12}$$\pm 0.08$
Refined $\mathbf{34.04}$$\pm 0.19$ $\mathbf{14.59}$$\pm 0.08$ $7.88$$\pm 0.04$ $6.24$$\pm 0.06$
@@ 3.4 Large Language Model size ablations
Figure 5: Normalized learning rate schedules learned on vanilla Transformer-based LLM trained on C4 dataset. Left: schedules obtained when setting $w_{t}\propto\|g_{t}\|^{-2}$. Right: schedules when $w_{t}\propto 1/\|g_{t}\|_{1}$. We find that $\ell_{1}$ norm is much more consistent across model sizes and baseline schedules. Thus, we used $w_{t}\propto 1/\|g_{t}\|_{1}$ weighting from linear baselines to obtained Refined schedule mentioned in Table 5. The x axis is training steps.
Table 5: LLM Ablations (C4 validation loss).
Schedule 117M 284M 1B 3.5B
Cosine $3.089$ $2.891$ $2.729$ $2.631$
Linear Decay $3.087$ $2.888$ $2.725$ $\mathbf{2.625}$
Refined $\mathbf{3.075}$ $\mathbf{2.884}$ $\mathbf{2.722}$ $2.634$
In addition to the results above, we validate our insights on Language Models by performing an additional set of experiments directly comparing to the setting explored in Chinchilla (Hoffmann et al., 2022) where we train a vanilla Transformer model on the C4 dataset (Raffel et al., 2020) for 10k steps with a token batch size of around 524k ($2^{19}$). Starting with Hoffmann et al. (2022), most recent LLMs employ the AdamW optimizer with Linear Warmup and Cosine Decay. We perform a head-to-head comparison of Cosine decay, Linear decay and Refined schedules (the latter is refined from Linear decay). As shown in Table 5, contrary to popular wisdom, Linear decay performs better than Cosine decay across all model sizes. The Refined schedule outperforms Linear Decay for all but the 3.5B sized model.
3.5 Limitations
The cifar10 and cifar100 problems show potential pitfalls from using our approach (Figure 4) when our assumptions break down. On these problems $100\%$ train accuracy is reached near the end, driving the gradient norm sequence to zero. The resulting refined schedule rapidly increases the step size near the end of optimization, which if used results in training divergence. We suggest using a linear decay schedule instead on problems where the gradient norm drops significantly at the end.
@@ 4 Discussion
Schedules of this form are already in use, with $p$ typically tuned as a hyper-parameter. From Figure 1, we see that $p<1$ should be used when the gradient norm is decreasing over time, $p>1$ when the gradient norm is increasing, and $p=1$ when the gradient sequence is flat. We use the term polynomial decay for the schedule given in Equation 3, which its name in both PyTorch and Tensorflow, although in the literature schedules of the form $1/t^{\alpha}$ are also sometimes referred to as polynomial schedules. From Figure 2, we can see that many of our test problems have refined schedules that are well-approximated by a polynomial decay sequences, shown in blue.
@@ 5 Related work
The Robbins-Monro conditions are the foundation of early learning rate theory (Robbins and Monro, 1951). They advocated for step size sequences where $\sum_{t=1}^{\infty}\eta_{t}=\infty$ and $\sum_{t=1}^{\infty}\eta_{t}^{2}<\infty$. Of the schedules satisfying these conditions, they advocated for schedules with $1/t$ decay as they are asymptotically optimal for twice continuously differentiable and strongly convex functions. This schedule was later shown to be optimal (even non-asymptotically) for strongly convex G-Lipschitz stochastic optimization when appropriate averaging is used (Shamir and Zhang, 2013; Lacoste-Julien et al., 2012; Rakhlin et al., 2012).
Nemirovski et al. (2009) point out that the performance of these step-sizes is extremely poor when the base step size is miss-specified, and advocate $1/\sqrt{t}$ step sizes, which have the advantage of performing asymptotically optimally when strong convexity is not present. The use of $1/\sqrt{t}$ step sizes in combination with averaging was originally developed by Polyak (1990) and Ruppert (1988), and are optimal for the complexity class of G-Lipschitz convex functions with or without stochasticity (Blair, 1985; Nesterov, 2018).
Scheduling for the stochastic quadratic setting has seen significant recent interest. The seminal work of Bach and Moulines (2013) establish that a flat $1/4R$ schedule (for a given problem dependent-constant $R$), independent of the stopping time $T$, gives the optimal $\mathcal{O}(1/T)$ rate for the average iterate. For the last iterate, Ge et al. (2019a) establish that the step-decay schedule’s rate:
$\Theta\left(\frac{d\sigma^{2}}{T}\log T\right),$
Table 6: Polynomial Decay Schedules fit to refined schedules
Problem Warmup Power
DLRM 20% 0.9
GPT 28% 1.0
RoBERTa 45% 0.8
MRI 40% 0.9
RCNN 0% 1.5
ImageNet 1% 3.0
is a multiplicative log factor worse than the mini-max optimal rate. Here $d$ is the dimensionality of the problem and $\sigma$ a bound on the noise. They further show that the any-time last-iterate behavior of SGD is poor, visiting highly sub-optimal function values infinitely often. Pan et al. (2022) study the situation where the full Hessian spectrum is know in advance. In this case, they are able to improve the rate to
$\mathcal{O}\left(\frac{d\sigma^{2}}{T}\log\kappa\right),$
or, under a skewed spectrum assumption, to
$\mathcal{O}\left(\frac{d\sigma^{2}}{T}\right).$
Ge et al. (2019b) show that commonly used schedules such as polynomial decay schedules are sub-optimal in the quadratic setting.
5.1 Non-smooth last iterate convergence
For non-smooth stochastic optimization, Shamir and Zhang (2013) show that the last iterate of SGD on non-smooth, convex problems with a $1/\sqrt{t}$ step size is
$\mathcal{O}\left(\frac{\log(n)}{\sqrt{n}}\right).$
Harvey et al. (2019) show a corresponding lower bound, showing that the log factor is unavoidable when using $1/\sqrt{t}$ step-sizes.
Jain et al. (2019) were the first to show that the log-factor can be removed if the horizon is known in advance. Their approach uses a sequence of step sizes:
$\gamma_{t}=\frac{D2^{-i}}{G\sqrt{T+1}},$
where $i$ is an epoch counter, for a sequence of epochs each half the length of the preceding epoch. This step size sequence gives a convergence rate for the last iterate of:
$E\left[f(x_{T})-f_{*}\right]\leq\frac{15GD}{\sqrt{T}}.$
For more general conversions from regret bounds to last iterate guarantees, Cutkosky (2019) provides an alternative conversion based upon evaluating gradients at running averages of iterates and achieves the same last-iterate convergence rate as Theorem 1. However, our method is directly interpretable in terms of learning rate schedules to yield immediate practical guidance while the technique of Cutkosky (2019) is a more opaque transformation of the base algorithm. Nevertheless, there are interesting contrasts: the result of Cutkosky (2019) is ”anytime” in the sense that each individual iterate comes with an optimal convergence guarantee. In our case, only the last iterate has an optimal guarantee (in fact, Zamani and Glineur (2023) show that this restriction is necessary for any scheme based on learning rate schedules). Understanding the differences in these approaches is a valuable future direction.
5.2 Schedules in Applications
@@
Learning rate schedules used in applications usually fall within a few standard classes supported by major frameworks. To survey the use of schedulers in open source software, we performed a GitHub search on each of the schedulers in the PyTorch Library. Table 7 shows that the polynomial decay schedule that we advocate is currently the least popular scheduler in PyTorch.
PyTorch Scheduler Github Files (in thousands)
ReduceLROnPlateau 105
StepLR 101
MultiStepLR 37.9
CosineAnnealingLR 37.1
ExponentialLR 16
OneCycleLR 14.9
CosineAnnealingWarmRestarts 10.9
CyclicLR 9.1
LinearLR 5.9
ConstantLR 3.6
MultiplicativeLR 2.6
PolynomialLR 1.3
Table 7: Popularity of standard learning rate schedules
The reduce-on-plateau, and its fixed length sister, the step-wise schedule, popularized by Krizhevsky et al. (2012) for use in image classification, is considered a benchmark method. They justify it heuristically (“The heuristic which we followed was to divide the learning rate by 10 when the validation error rate stopped improving with the current learning rate.”). Due to its use in highly influential papers as well as its use in tutorials, it is by far the most commonly used schedule (Table 7).
Many early works use step-sizes satisfying the Robbins-Monro conditions, such as step sizes proportional to $1/(c+\alpha t)$, for some $c$ and $\alpha$, as advocated in the influential work of Bottou (1991, 2012). These step sizes produce large values for small $t$ and so are often clipped (using $\min(\beta,\cdot)$ for some $\beta$).
Schedules involving restarting (also called cycling) are also widely used, particularly on computer vision problems where overfitting can be an issue (Smith and Topin, 2018). Loshchilov and Hutter (2017b) use restarting together with the cosine annealing schedule:
$\eta_{t}=\eta_{\min}+\left(\eta_{\max}-\eta_{\min}\right)\left(1+\cos\left(\frac{t}{T}\pi\right)\right),$ (4)
This schedule is remarkably effective in practice for image classification problems even without restarts; as far as the authors are aware, for ResNet-50 ImageNet training, no schedule has been demonstrated that out-performs it. It is less commonly used in other problem domains. The motivation for the use of cosine function is unstated, and there is no theory to support it. The reasons for the success of the schedule are not currently understood.
The only large-scale comparison of schedulers for deep learning optimization that we are aware of is the work of Schmidt et al. (2021). They compare on a smaller set of problems than our comparison, although across a range of optimizers, whereas we focus on just Adam & SGD.
5.3 Any-time Adaptive Schedules
@@
There has been significant work on any-time learning rate adaptation that makes use of gradient or loss information to adapt the learning rate. Schedules in the AdaGradNorm class are backed by strong theory (Wu et al., 2020; Streeter and McMahan, 2010; Duchi et al., 2011; Ward et al., 2019; Li and Orabona, 2019). They use step sizes of the form:
$\eta_{t}\propto\frac{1}{\sqrt{\sum_{t=1}^{T}\left\|g_{t}\right\|^{2}}}$
They have not seen significant adoption in practice, likely due to the intrinsic disadvantages of any-time $O(1/\sqrt{t})$-style schedules as already discussed. It is also not clear how to apply this schedule on top of existing high-performance optimizers such as Adam.

### v2

@@ 3 Experiments
For each problem considered, we first performed a sweep of learning rates on a grid [$10^{i}$, $2\times 10^{i}$, $5\times 10^{i}$] for varying $i$, separately for each schedule. We then ran multiple seeds using the best learning-rate from the grid search. Mean and standard error of the mean was tabulated for each schedule. The best result for each method, up to a statistical significance level of 0.05 using a paired two-sample t-test, is highlighted in bold. Specific details of the hyper-parameters used are given in Appendix G. All non-adaptive schedules included a fixed learning-rate warmup, with length following standard practices for the problem. Our step-wise schedule divides by 10 at 30-60-90 percent completion, following standard practices from ImageNet training. The Refined schedules were calculated using the linear schedule to generate the initial gradient norm sequence.
3.1 Convex problems
To validate the effectiveness of our method on convex problems, we performed a comparison across 8 commonly used classification problems from the LIBSVM repository, with separable problems excluded (see Section A). We used a logistic regression loss together with the Adam optimizer with $\beta=(0.9,0.95)$, with batch size 16 and trained for 100 epochs. Table 1 demonstrates that both the linear decay schedule and our refinement schedule consistently either match or out-perform the cosine schedule. The linear decay schedule matches the cosine schedule on every problem (up to statistical significance), and out-performs it on two problems. Overall, the Refined schedule out-performs all previous baseline schedules on the majority of problems.
Figure 3: Gradient Norm sequences and the resulting Refined schedules, generated using an initial linear decay schedule with warmup for the initial run. Log scale views are inset for scale reference.
@@ 3.3 Large-scale schedule benchmark
We performed a large-scale comparison of schedules across common deep learning benchmarks, (pre-)training each problem from scratch for each task:
@@
For a small-scale image-classification problem, we chose CIFAR10 (Krizhevsky,, 2009). A high-performance Wide ResNet architecture was used (Zagoruyko and Komodakis,, 2016). Test error percentage is reported.
@@
A more realistic yet still small scale benchmark. We used a compact DenseNet architecture (Huang et al.,, 2017) to increase the diversity of model architectures tested. Test error percentage is reported.
@@
We used a ResNet-50 architecture (He et al.,, 2016) for the ImageNet (Russakovsky et al.,, 2015) classification task. We use a standard data-augmentation pipeline following other recent comparisons in the optimization literature. Test error percentage is reported.
@@
A small-scale German-English translation task (Cettolo et al.,, 2014) using a LSTM model (Wiseman and Rush,, 2016). Test Perplexity is reported.
@@
A modern small (162M parameter) GPT-style auto-regressive transformer model (Radford et al.,, 2019) trained on the large Book-Wiki corpus (Zhu et al.,, 2015). Test Perplexity is reported.
@@
A masked autoencoder variant (Liu et al.,, 2019) also trained on the Book-Wiki corpus. Test Perplexity is reported.
@@
A high-performance ImageNet classifier using a Vision Transformer (Dosovitskiy et al.,, 2021). Test error percentage is reported.
@@
A modern recommendation system engine (Naumov et al.,, 2019) trained on the open Criteo Kaggle Display Advertising dataset. Test Accuracy is reported.
@@
A large stacked U-Net architecture (Sriram et al.,, 2020) trained on the fastMRI dataset (Zbontar et al.,, 2018), an image-to-image regression task from the medical imaging domain. Test set metric $100\cdot(1-\text{SSIM})$ is reported.
@@
Figure 4: Normalized learning rate schedules learned on vanilla Transformer-based LLM trained on C4 dataset. Left: schedules obtained when setting $w_{t}\propto\|g_{t}\|^{-2}$. Right: schedules when $w_{t}\propto 1/\|g_{t}\|_{1}$. We find that $\ell_{1}$ norm is much more consistent across model sizes and baseline schedules. Thus, we used $w_{t}\propto 1/\|g_{t}\|_{1}$ weighting from linear baselines to obtained Refined schedule mentioned in Table 4(b). The x axis is training steps.
@@
Problem Flat 1/t 1/sqrt Offset 1/t Offset 1/sqrt Linear CIFAR10 $8.04$$\pm.13$ $5.42$$\pm.28$ $6.37$$\pm.41$ $9.23$$\pm.08$ $6.58$$\pm.06$ $\mathbf{4.35}$.$\pm.05$ CIFAR100 $30.43$$\pm.20$ $26.58$$\pm.11$ $29.09$$\pm.16$ $32.80$$\pm.07$ $27.62$$\pm.13$ $\mathbf{22.11}$.$\pm.08$ ImageNet $33.00$$\pm.14$ $26.48$$\pm.06$ $28.35$$\pm.05$ $47.94$$\pm.08$ $27.34$$\pm.06$ $\mathbf{23.11}$.$\pm.07$ IWSLT14 $8.07$$\pm.02$ $7.62$$\pm.01$ $7.52$$\pm.01$ $12.89$$\pm.06$ $8.48$$\pm.01$ $\mathbf{7.10}$.$\pm.01$ GPT $20.20$$\pm.000$ $18.99$$\pm.04$ $19.48$$\pm.02$ $27.85$$\pm.05$ $22.88$$\pm.006$ $\mathbf{18.60}$.$\pm.02$ RoBERTa $4.52$$\pm.005$ $4.25$$\pm.007$ $4.33$$\pm.01$ $5.33$$\pm.02$ $5.15$$\pm.02$ $\mathbf{3.94}$.$\pm.007$ DLRM $\mathbf{20.95}$.$\pm.006$ $47.59$$\pm 6.45$ $45.99$$\pm 5.98$ $\mathbf{20.94}$.$\pm.007$ $20.99$$\pm.009$ $\mathbf{20.94}$.$\pm.006$ MRI $\mathbf{9.00}$.$\pm.04$ $8.91$$\pm.01$ $\mathbf{8.98}$.$\pm.04$ $9.53$$\pm.08$ $9.16$$\pm.05$ $\mathbf{8.88}$.$\pm.02$ ViT $30.11$$\pm.27$ $28.36$$\pm.40$ $28.53$$\pm.15$ $73.84$$\pm 6.08$ $50.36$$\pm 12.39$ $\mathbf{24.82}$.$\pm.31$ RCNN $65.43$$\pm.12$ $63.38$$\pm.05$ $64.13$$\pm.10$ $79.32$$\pm.07$ $69.25$$\pm.07$ $\mathbf{60.98}$.$\pm.02$
@@
RCNN $61.76$$\pm.06$ $\mathbf{61.00}$.$\pm.04$ $\mathbf{60.98}$.$\pm.02$ $61.21$$\pm.03$
Schedule 1 Epoch 5 Epochs 30 Epochs Cosine $35.80$$\pm 0.27$ $15.07$$\pm 0.11$ $\mathbf{6.08}$$\pm 0.03$ Lin. Decay $\mathbf{33.84}$$\pm 0.17$ $\mathbf{14.36}$$\pm 0.07$ $\mathbf{6.12}$$\pm 0.08$ Refined $\mathbf{34.04}$$\pm 0.19$ $\mathbf{14.59}$$\pm 0.08$ $6.24$$\pm 0.06$
(a) CIFAR10 Training Time Ablations (Test Error %)
Schedule 117M 284M 1B 3.5B Cosine $3.089$ $2.891$ $2.729$ $2.631$ Linear Decay $3.087$ $2.888$ $2.725$ $\mathbf{2.625}$ Refined $\mathbf{3.075}$ $\mathbf{2.884}$ $\mathbf{2.722}$ $2.634$
(b) LLM Ablations (C4 validation loss)
Tables 2 & 3 show the results. We break the schedules into two categories, classical and modern. The modern schedules consistently outperform the classical schedules, often by large margins. Although this is common folk-law in deep learning, we are not aware of any existing large-scale experiments establishing this. Our comparison of modern schedules shows a clear hierarchy among the schedules. The Stepwise schedule is dominated by the Cosine schedule, and the Linear schedule matches or outperforms the Cosine schedule on all problems except ViT. The refined schedule further outperforms the Linear schedule on 5 of the 10 problems, but shows mixed results on ViT and RCNN. The refinement process produced degenerate schedules that fail on the CIFAR problems, we discuss this in Appendix A.
@@ 3.4 Cosine schedule ablations
To further investigate the failure modes of the cosine schedule, we ran a series of shorter duration training runs on CIFAR10. Our premise was that the cosine schedule is heavily over-fit to long training duration computer vision problems. As shown in Table 4(a), the cosine schedule begins to under-perform the linear decay schedule and the refined schedules when training for less than 30 epochs. In contrast, while the refined schedule also under-performs for longer duration training it has no statistically significant differences from the linear decay schedule for shorter duration training, where they both perform consistently better than the cosine schedule.
@@ 3.5 Large Language Model size ablations
In addition to the results above, we validate our insights on Language Models by performing an additional set of experiments directly comparing to the setting explored in Chinchilla (Hoffmann et al.,, 2022) where we train a vanilla Transformer model on the C4 dataset (Raffel et al.,, 2020) for 10k steps with a token batch size of around 524k ($2^{19}$) and learning rate $0.0002$. Starting with Hoffmann et al., (2022), most recent LLMs employ the AdamW optimizer with Linear Warmup and Cosine Decay. We perform a head-to-head comparison of Cosine decay, Linear decay and Refined schedules (the latter is refined from Linear decay). As shown in Table 4(b), contrary to popular wisdom, linear decay performs better than cosine decay across all model sizes. The Refined schedule further outperforms linear decay for all but the 3.5B sized model.
@@ 4 Discussion
Although our theoretical analysis assumes that the underlying optimization problems are convex, as we have demonstrated, our method appears to work well for non-convex deep learning problems in practice. A similar theoretical analysis in the non-convex setting can not be done as last iterate rates are provably impossible in that setting (Drori and Shamir,, 2020).
@@
Schedules of this form are already in use, with $p$ typically tuned as a hyper-parameter. From Figure 1, we see that $p<1$ should be used when the gradient norm is decreasing over time, $p>1$ when the gradient norm is increasing, and $p=1$ when the gradient sequence is flat. We use the term polynomial decay for the schedule given in Equation 2, which its name in both PyTorch and Tensorflow, although in the literature schedules of the form $1/t^{\alpha}$ are also sometimes referred to as polynomial schedules.
@@ 5 Related Work
The standard schedule for strongly-convex stochastic optimization is $\eta_{t}\propto 1/t$. This is optimal when appropriate averaging is used (Shamir and Zhang,, 2013; Lacoste-Julien et al.,, 2012; Rakhlin et al.,, 2012). Without strongly convexity, theory typically recommends a constant schedule.
@@
Learning rate schedules used in applications usually fall within a few standard classes supported by major frameworks. To survey the use of schedulers in open source software, we performed a GitHub search on each of the schedulers in the PyTorch Library. Table H (Appendix) shows that the polynomial decay schedule, which includes the linear decay schedule as a special case, is currently the least popular scheduler in PyTorch. Step-wise schedules are by far the most popular.
Our approach is motivated by the remarkable work of Pan et al., (2022), who show that for the last iterate of quadratic problems, minimizing an upper bound can yield improved schedules. However their work requires knowledge of the full Eigenspectrum of the Hessian, making it impractical to use. Our theory relies the sequence of expected gradient norms, a more tractable quantity.
Ours is not the first reduction from regret bounds to a last-iterate guarantee: Cutkosky, (2019) suggests evaluating gradients at running averages to achieve the same convergence rate as Theorem 1. However, our method is directly interpretable in terms of learning rate schedules to yield immediate practical guidance while the previous technique is a more opaque transformation of the base algorithm. Nevertheless, understanding the differences is a valuable future direction.
@@
There has been significant work on learning rate adaptation using gradient or loss information. The AdaGradNorm-type schedule is backed by strong theory (Wu et al.,, 2020; Streeter and McMahan,, 2010; Duchi et al.,, 2011; Ward et al.,, 2019; Li and Orabona,, 2019). They suggest $\eta_{t}\propto 1/\sqrt{\sum_{t=1}^{T}\left\|g_{t}\right\|^{2}}$, but it is unclear how to apply these schedules to other optimizers such as Adam.